# The Structure of the Majorana Clifford Group

### v2

@@ II.5 Hilbert Space Representation
II.6 Majorana Algebras
While taking the trace of a Majorana string (or any operator that can be expanded in terms of Majorana strings) is straightforward, the lack of a tensor product structure in (3) means that the Hilbert space has no obvious factorization with regard to the individual mode operators $\chi_{i}$, and hence no way to take a partial trace. Instead one has to adopt the notion of Majorana (sub-) algebras [25], which we briefly review here due to their importance e.g. for analyzing the entanglement and entropy of fermionic subsystems.
Any operator $O$ acting on the Hilbert space $\operatorname{\mathcal{H}}$ that can be expressed as a linear combination of Majorana strings such that
$O\equiv\sum_{v\in\mathbb{F}_{2}^{2n}}c_{v}\cdot\operatorname{\mu}(v),\quad c_{v}\in\mathbb{C},$ (34)
is an element of the (fundamental) Majorana algebra. In fact, it is straightforward to show that a Majorana algebra composed of $2n$ modes encompasses all complex-valued $2^{n}\times 2^{n}$ matrices. Using (32) and the fact that Hermitean strings square to the identity, the coefficients $c_{v}$ of the expansion can be determined using:
$c_{v}=\frac{1}{2^{n}}\operatorname{tr}(O\operatorname{\mu}(v)).$ (35)
where $n_{A}=|A|/2$, and $\Pi$ is some fixed permutation matrix. Note however that we run into issues with regard to the Hilbert space representation of the subalgebra. The strings as depicted in (36) are still endowed with the ordinary $2^{n}\times 2^{n}$-representation (32), even though the algebra itself is isomorphic to the $2^{n_{A}}\times 2^{n_{A}}$ algebra of complex matrices. Hence this representation is reducible, which is especially problematic when considering properties like entropy that intrinsically depend on the Hilbert space dimension. To compensate, we introduce a new irreducible representation $\operatorname{\mu}_{A}(v_{A})$ which satisfies (3) and (32) with $n\rightarrow n_{A}$. The reduced operator $O_{A}$ corresponding to the chosen subalgebra is then given by
III The Fermionic Clifford Group
@@
The Majorana Clifford group $\operatorname{\mathcal{C}}_{2n}$ is the group of all unitary automorphisms acting on the Majorana group $\operatorname{\mathcal{M}}_{2n}$ (5), i.e. all unitary operators $\operatorname{\gamma}$ that map (general) Majorana strings to other such strings:
@@
for some functions $c:\mathbb{F}_{2}^{2n}\rightarrow\operatorname{\mathbb{C}}$ and $S:\mathbb{F}_{2}^{2n}\rightarrow\mathbb{F}_{2}^{2n}$ that depend on $\operatorname{\gamma}$.
@@ III.1 Symplectic Structure
The same must therefore be true for elements of the Clifford group. Since it was shown in section II.2 that two Majorana strings $\operatorname{\mu}(v),\operatorname{\mu}(v^{\prime})$ (anti-) commuting is equivalent to the symplectic product $\left\langle v,v^{\prime}\right\rangle$ being zero (or one), the function $S$ associated to a Clifford element as described in definition 5 has to preserve the structure of said product. We later prove that any $S$ that does so is necessarily linear and hence symplectic:
@@
Obtaining the associated group element in the Pauli basis (where one would require $S^{T}\eta S=\eta$ instead) is then as simple as performing the Jordan-Wigner transformation (30):
@@
A proof for this discrete variation of the Stone-von Neumann Theorem is found in Appendix A.1.
Note that it is not possible to find a closed form expression for the coefficients $c(v)$ in (38). This is due to the field $\mathbb{F}_{2}$, which describes Majorana (and qubit Pauli) strings, being special from a mathematical perspective. Still, by evaluating $\operatorname{\gamma}(S)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\gamma}^{\dagger}(S)$ in two different ways one can obtain the following relation:
$c(v)\,c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(Sv,Sv^{\prime})}\cdot c(v+v^{\prime}),$ (46)
which implies that $c(v)^{2}=c(0)=1$ due to (12), as expected from unitarity. Stronger statements can be made for qudit Pauli strings with $d>2$ prime [23], but since the overall phase of a quantum state has no physical meaning, we can safely ignore it most of the time, and focus solely on the symplectic representation $S$. $\operatorname{\gamma}(S)$ is therefore used to denote any Clifford that acts accordingly.
@@ III.2 Parity-Preserving Cliffords
So far the Pauli and Majorana Clifford groups are mathematically equivalent. But unlike Pauli strings, Majorana strings have the additional extrinsic property of fermion parity. It is easy to see that Clifford operators do not obey parity superselection in general, e.g.
@@
depicts a symplectic matrix (Clifford operator) mapping between vectors (Majorana strings) of different parity. For an element $\operatorname{\gamma}(S)$ to respect the parity of any string, the associated symplectic $S$ must satisfy
@@
Overall, we can therefore identify a subgroup of the Clifford group $\operatorname{\mathcal{C}}_{2n}$ that acts according to $\operatorname{O}(2n)$ and hence preserves parity:
@@
Note that we have $\operatorname{\mathcal{M}}_{2n}^{+}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, because while odd-parity strings don’t affect the parity of other strings when acting from both sides, they can indeed change the parity of states in the Hilbert space. This is shown explicitly in Section IV.2. Overall, all elements of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ hence commute with $(-1)^{F}$ (19) because
@@
where $Sj=j$ due to all $S\in\operatorname{O}(2n)$ having odd-parity rows, and $c(j)=1$ using arguments made in Appendix A.2.
III.3 Generating Set
For the ordinary Pauli Clifford group, the Gottesman-Knill theorem [28] states (among other things) that $\operatorname{\mathcal{C}}_{2n}$ can be efficiently generated from repeated applications of tensor-products containing CNOT, Hadamard, and the phase gate (or an equivalent minimal set) as the only non-trivial operators. In the symplectic representation, this is equivalent to the fact that any element $S\in\operatorname{Sp}(2n)$ can be decomposed into a product of symplectic transvections
$\tau_{a}:v\mapsto v+\left\langle a,v\right\rangle_{\operatorname{\pi}}a,\quad a\in\mathbb{F}_{2}^{2n}.$ (57)
Additionally, all transvections can be written as products and (in the Pauli basis [^5] In the Majorana basis it is straightforward to show that the direct sum of symplectics is not necessarily symplectic anymore. More specifically, only direct sums of orthogonal matrices are again orthogonal and therefore symplectic.) direct sums of just 3 independent ones:
  • For $n=1$, there are two such independent transvections (e.g. for $s=(1,0)^{T}$ and $h=(1,1)^{T}$), from which all other symplectics can be generated by repeated multiplication of $\tau_{s}$ (which corresponds to the phase gate) and $\tau_{h}$ (which corresponds to the Hadamard).
  • For $n=2$, it can be shown that there is only one independent transvection $\tau_{c}$ (for e.g. $c=(1,0,0,1)^{T}$) since any other one can be written as $\tau_{c}$ times some direct sum of $n=1$ symplectics. In fact, CNOT can be shown to correspond to the symplectic $(\tau_{s}\oplus\tau_{h}\tau_{s}\tau_{h})\,\tau_{c}$ in the Pauli basis.
@@ III.3 Generating Set
We now present an analogous statement for just the p-Clifford subgroup $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$. Our focus hereby is only on generating the orthogonal action since adding a phase change-inducing element of $\operatorname{\mathcal{M}}^{+}_{2n}$ is trivial.
@@
Any element of the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}\leq\operatorname{\mathcal{C}}_{2n}$ (mod $\operatorname{\mathcal{M}}^{+}_{2n}$, and excluding $\operatorname{BRAID}_{2n}(j)$ due to (56)) can be efficiently generated from products of $\operatorname{BRAID}_{2}$ and $\operatorname{BRAID}_{4}$ operators, which for any $k\geq 0$ are defined as:
$\begin{split}\operatorname{BRAID}_{k}(a)&\coloneqq\exp\left(i\frac{\pi}{4}\operatorname{\mu}(a)\right)\\ &=\frac{1}{\sqrt{2}}\big{(}I+i\operatorname{\mu}(a)\big{)},\end{split}$ (58)
with $a\in\mathbb{F}_{2}^{2n},\,\operatorname{w}(a)=k$ (compare to [13, 15]).
@@
We first note the well-known fact that the orthogonal group $\operatorname{O}(2n)$ is generated by the set of Householder reflections [26]
$h_{a}:v\mapsto v+(a^{T}v)\,a,\quad a\in\mathbb{F}_{2}^{2n,+}.$ (59)
Compare this to the action of $\operatorname{BRAID}_{k}(a)$ on a general Majorana string $\operatorname{\mu}(v)$, which can be shown to be
$\begin{split}&\operatorname{BRAID}_{k}(a)\operatorname{\mu}(v)\operatorname{BRAID}^{\dagger}_{k}(a)\\ ={}&\frac{1}{2}\left(1+(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(v)\\ &+\frac{i}{2}\left(1-(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(a)\operatorname{\mu}(v)\\ \equiv{}&(i\,\zeta(a,v))^{a^{T}v}\cdot\operatorname{\mu}(v+(a^{T}v)\,a),\end{split}$ (60)
using the fact that $\left\langle a,v\right\rangle=a^{T}v$ for $\operatorname{p}(a)=0$ and any $v$. The symplectic representations of the (even-parity) $\operatorname{BRAID}_{k}$ operators are hence the Householder reflections:
$\operatorname{BRAID}_{k}(a)\equiv\operatorname{\gamma}(h_{a}),$ (61)
which implies that they generate the orthogonal action of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ [^6] This argument can be extended to the whole Clifford group by also including braiding operators with odd weights. In fact, symplectic transvections reduce to Householder reflections when restricting $a$ in (57) to be even.. In appendix B we provide an algorithm that can generate any orthogonal matrix from Householder reflections in $\mathcal{O}(n^{3})$ steps.
What is left to show is that all reflections (except $h_{j}$) can be generated from just the ones corresponding to strings with weight 2 and weight 4. This can be done in at most $\mathcal{O}(n^{3})$ steps by noting that one can write $h_{a}=I+aa^{T}$, and hence
@@
Given some reflection associated to a vector $a$ with weight $k>4$ even, the problem therefore reduces to finding weight 4 reflections that map it to a vector with weight 2 or 4. This is always possible because any weight 4 Householder is of the form $\operatorname{\omega}_{4}\oplus I_{2n-4}$ (up to permutations), essentially allowing us to use (20) with regard to any 4 (odd-parity) entries of $a$ to reduce its weight by 2. This reduction has to be performed $\mathcal{O}(k)$ times, each step involving two $\mathcal{O}(n^{2})$ matrix applications due to (62). ∎
@@ IV Fermionic Stabilizers
We don’t mention explicit examples of fermionic stabilizer codes as they can be readily found in e.g. [9, 15, 12, 10, 11].
@@ IV.1 Symplectic Representation
The reason why a stabilizer can have at most $n$ independent generators becomes apparent when one considers the symplectic equivalent:
@@
Due to fermion superselection we only care about isotropic subspaces of the even-parity subspace $\mathbb{F}_{2}^{2n,+}$. In this case each maximally isotropic subspace if of the form
@@
where $M_{\text{std}}$ is the canonical maximally isotropic subspace given in terms of its basis as row vectors of a $n\times 2n$ matrix
@@
where $M\subset\mathbb{F}_{2}^{2n,+}$ is an even-parity isotropic subspace of dimension $r=n-k$, and $v$ is an element of the quotient set $\mathbb{F}_{2}^{2n}/M$  [^7] A priori, we allow $v$ to have odd parity, as it only affects the leading phases. We will return to this in section IV.2..

### v3

@@ II.5 Hilbert Space Representation
While taking the trace of a Majorana string (or any operator that can be expanded in terms of Majorana strings) is straightforward, the lack of a tensor product structure in (4) means that the Hilbert space has no obvious factorization with regard to the individual mode operators $\chi_{i}$, and hence no way to take a partial trace. Instead one has to adopt the notion of Majorana (sub-) algebras [28]. We provide the basics of this formalism in Appendix B due to its importance e.g. for analyzing the entanglement and entropy of fermionic subsystems.
III The Majorana Clifford Group
@@
The Majorana Clifford group $\operatorname{\mathcal{C}}_{2n}$ is the group of all unitary automorphisms acting on the Majorana group $\operatorname{\mathcal{M}}_{2n}$ (6), i.e. all unitary operators $\operatorname{\Gamma}$ that map (general) Majorana strings to other such strings (up to a phase):
@@
for some functions $c:\mathbb{F}_{2}^{2n}\rightarrow U(1)$ and $S:\mathbb{F}_{2}^{2n}\rightarrow\mathbb{F}_{2}^{2n}$ that depend on $\operatorname{\Gamma}$.
@@ III.1 Symplectic Structure
The same must therefore be true for elements of the Clifford group. Since it was shown in section II.2 that two Majorana strings $\operatorname{\mu}(v),\operatorname{\mu}(v^{\prime})$ (anti-) commuting is equivalent to the symplectic product $\left\langle v,v^{\prime}\right\rangle$ being zero (or one), the function $S$ associated to a Clifford element as described in definition 5 has to preserve the structure of said product. We later prove in Theorem 1 that any $S$ that does so is necessarily linear and hence symplectic:
@@
Obtaining the associated group element in the Pauli basis (where one would require $S^{T}\eta S=\eta$ instead) is then as simple as applying a Jordan-Wigner transformation (29):
@@
A proof for this discrete variation of the Stone-von Neumann Theorem [^4] We call it a Stone-von Neumann theorem in accordance with the equivalent theorem for Pauli Cliffords [23]. There one can establish an explicit connection between the commutation relations of $Z$ and $X$, and the commutation relations of position and momentum operators. is found in Appendix D.1.
Note that it is not possible to find a closed form expression for the coefficients $c(v)$ in (36). This is again due to the nontrivial composition rules (7) of two Majorana strings. Still, by evaluating $\operatorname{\Gamma}(S)\operatorname{\mu}(v)\operatorname{\mu}(v^{\prime})\operatorname{\Gamma}^{\dagger}(S)$ in two different ways one can obtain the following relation:
$c(v)\cdot c(v^{\prime})=\frac{\zeta(v,v^{\prime})}{\zeta(Sv,Sv^{\prime})}\cdot c(v+v^{\prime}),$ (44)
which implies that $c(v)^{2}=c(0)=1$ due to (13) and $\operatorname{\mu}(0)=I$. Hence one has
$c(v)=\pm 1\quad\forall v\in\mathbb{F}_{2}^{2n},$ (45)
as expected from unitarity. Stronger statements can be made for qudit Pauli strings with $d>2$ prime [23], but since the overall phase of a quantum state has no physical meaning [^5] The phases do however become important when considering superpositions of Majorana strings or quantum states, in which case a more detailed analysis is necessary. This could be explored in future work., we can safely ignore it most of the time and focus solely on the symplectic representation $S$. $\operatorname{\Gamma}(S)$ is therefore used to denote any Clifford unitary that acts accordingly.
@@ III.2 Parity-Preserving Cliffords
So far the Pauli and Majorana Clifford groups are mathematically equivalent. But unlike Pauli strings, Majorana strings have the additional extrinsic property of operator parity. It is easy to see that Clifford operators do not preserve thar parity in general, e.g.
@@
depicts a symplectic matrix (Clifford operator) mapping between vectors (Majorana strings) of different parity. This can be interpreted as observables not being mapped to other observables under such transformations, which is not physically reasonable.
For an element $\operatorname{\Gamma}(S)$ to respect the parity of any string, the associated symplectic $S$ must satisfy
@@
We can therefore identify a subgroup of the Clifford group $\operatorname{\mathcal{C}}_{2n}$ that acts according to $\operatorname{O}(2n)$ and thus preserves parity:
@@
Note that we have $\operatorname{\mathcal{M}}_{2n}^{+}\leq\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$, because while odd-parity strings don’t affect the operator parity of other strings when acting from both sides, they can indeed change the parity of states in the Hilbert space. This is shown explicitly in Section IV.2. Overall, all elements of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$ hence commute with $(-1)^{F}\equiv\operatorname{\mu}(j)$ (19) because
@@
where $Sj=j$ due to all $S\in\operatorname{O}(2n)$ having odd-parity rows, and $c(j)=1$ using arguments for fixed points of $S$ made in Appendix D.2.
III.3 Generating Set & Unitary Representation
In the Pauli basis, the Gottesman-Knill theorem [31] states that the Clifford group $\operatorname{\mathcal{C}}_{2n}$ can be efficiently generated from repeated applications of tensor-products containing CNOT, Hadamard, and the phase gate (or an equivalent minimal set) as the only non-trivial operators. In the symplectic representation, this is equivalent to the fact that any element $S\in\operatorname{Sp}(2n)$ can be decomposed into a product of symplectic transvections
$\begin{split}&\tau_{a}:v\mapsto v+\left\langle a,v\right\rangle_{\operatorname{\pi}}a\\ \Longleftrightarrow\quad&\tau_{a}=I+a\,a^{T}\operatorname{\eta},\quad\quad\quad a\in\mathbb{F}_{2}^{2n}.\end{split}$ (56)
Additionally, all such transvections can be written as products and (in the Pauli basis) direct sums of just 3 independent ones:
  • For $n=1$, there are two independent transvections (corresponding to e.g. $s=(0,1)^{T}$ and $h=(1,1)^{T}$), from which all other symplectics can be generated by repeated multiplication of $\tau_{s}$ (which represents the phase gate) and $\tau_{h}$ (which represents the Hadamard gate).
  • For $n=2$, it can be shown that there is only one independent transvection $\tau_{c}$ (for e.g. $c=(0,1,1,0)^{T}$) since any other one can be written as $\tau_{c}$ times some direct sum of $n=1$ symplectics. In fact, CNOT corresponds to the symplectic $(\tau_{s}\oplus\tau_{h}\tau_{s}\tau_{h})\,\tau_{c}$ in the Pauli basis.
@@ III.3 Generating Set & Unitary Representation
It is therefore reasonable to assume that similar statements also hold for $\operatorname{\mathcal{C}}_{2n}$ in the Majorana basis, and indeed there too the symplectic group is generated by transvections with regard to $\operatorname{\omega}$:
$\begin{split}W\tau_{a}W&=I+(Wa)(Wa)^{T}\operatorname{\omega}\\ &\equiv h_{(Wa)}.\end{split}$ (57)
The unitary representations of transvections in the Majorana basis are usually referred to as braiding operators, and can be defined as follows:
Definition 8 (Braiding Operators).
For any Hermitian Majorana string $\operatorname{\mu}(a)$ with $a\in\mathbb{F}_{2}^{2n}$, the associated braiding operator $\operatorname{BRAID}(a)\equiv\operatorname{BRAID}(\operatorname{\mu}(a))$ is defined as
$\begin{split}\operatorname{BRAID}(a)&\coloneqq\exp\left(i\frac{\pi}{4}\operatorname{\mu}(a)\right)\\ &=\frac{1}{\sqrt{2}}\big{(}I+i\operatorname{\mu}(a)\big{)}.\end{split}$ (58)
That these operators indeed act as symplectic transvections is straightforward to prove by letting them act on a Majorana string:
$\begin{split}&\operatorname{BRAID}(a)\operatorname{\mu}(v)\operatorname{BRAID}^{\dagger}(a)\\ ={}&\frac{1}{2}\left(1+(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(v)\\ &+\frac{i}{2}\left(1-(-1)^{\left\langle a,v\right\rangle}\right)\operatorname{\mu}(a)\operatorname{\mu}(v)\\ \equiv{}&(i\,\zeta(a,v))^{\left\langle a,v\right\rangle}\cdot\operatorname{\mu}(v+\left\langle a,v\right\rangle a),\end{split}$ (59)
where the last line can be derived by comparing the cases of $\left\langle a,v\right\rangle=0$ and $\left\langle a,v\right\rangle=1$. We can therefore identify
$\begin{split}\operatorname{\Gamma}(h_{a})&\sim\operatorname{BRAID}(a),\\ \quad c(v)&\sim(i\,\zeta(a,v))^{\left\langle a,v\right\rangle}\\ &=(-1)^{\left\langle a,v\right\rangle(a^{T}\operatorname{\operatorname{\omega}_{L}}v+f(a,v)+1)},\end{split}$ (60)
with the latter equality following from (8), thereby making the validity of (45) manifest. Note that $\operatorname{BRAID}^{\dagger}(a)$ acts in the exact same way, except for sign change according to $c(v)\rightarrow(-1)^{\left\langle a,v\right\rangle}c(v)$.
However, unlike in the Pauli case, direct sums of transvections in the Majorana basis (or equivalently fermionic tensor products of braiding operators) are not necessarily valid transvections anymore. In fact, it is only possible for transvections $h_{a}$ with $a\in\mathbb{F}_{2}^{2n,+}$ [^7] This can be shown e.g. by writing a generic direct sum of transvections as a block matrix and checking under which conditions it still satisfies (39)., in which case (20) dictates that they must be of the form
$\begin{split}&h_{a}:v\mapsto v+(a^{T}v)\,a\\ \Longleftrightarrow\quad&h_{a}=I+a\,a^{T},\quad\quad\quad a\in\mathbb{F}_{2}^{2n,+}.\end{split}$ (61)
Such transformations are called Householder reflections, and it should come as no surprise that they correspond to the parity-preserving subset of symplectic transvections. In the style of Gottesman-Knill, one can then prove that the corresponding set of braiding operators (efficiently) generates $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}/\operatorname{\mathcal{M}}^{+}_{2n}$:
@@
Any element of the p-Clifford group $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}\leq\operatorname{\mathcal{C}}_{2n}$ (mod $\operatorname{\mathcal{M}}^{+}_{2n}$) can be efficiently generated from products of braiding operators $\operatorname{BRAID}(a)$ with $a\in\mathbb{F}_{2}^{2n,+}$ as in (58). More specifically, it can be efficiently generated from the subset of braiding operators satisfying $\operatorname{w}(a)=2$, $\operatorname{w}(a)=4$ or $a=j$ (compare to [13, 15]).
@@
We first note the well-known fact that the (binary) orthogonal group $\operatorname{O}(2n)$ is generated by the set of Householder reflections as defined in (LABEL:eq:householder) [29]. This can be done in $\mathcal{O}(n^{3})$ steps, as demonstrated by the sampling algorithm we provide in appendix E. Using (LABEL:eq:braiding_action), it is then straightforward to argue that the Householder reflections efficiently generate the orthogonal action of $\operatorname{\mathcal{C}}^{\operatorname{p}}_{2n}$.
What is left to show is that all reflections (except $h_{j}\equiv\omega$) can be generated from just the ones corresponding to vectors with weight 2 and weight 4. This can be done in at most $\mathcal{O}(n^{3})$ steps by first noting that the Householder reflections exhibit a braiding algebra:
@@
Given some reflection associated to a vector $a$ with weight $k>4$ even, the problem therefore reduces to finding weight-4 reflections that map it to a vector with weight 2 or 4. This is always possible because any weight-4 Householder is of the form $\operatorname{\omega}_{4}\oplus I_{2n-4}$ up to permutations (which correspond to weight-2 reflections). This essentially allows us to use (20) with regard to any 4 entries of $a$ with local weight 3 to map them to their weight-1 complement, thereby reducing the overall weight by 2. For example, locally:
$\operatorname{\omega}h_{(0,1,1,1)^{T}}\operatorname{\omega}=h_{\operatorname{\omega}(0,1,1,1)^{T}}=h_{(1,0,0,0)^{T}}$ (63)
This reduction has to be performed $\mathcal{O}(k)$ times, each step involving two $\mathcal{O}(n^{2})$ matrix applications due to (62). ∎
Even though the nonlocality of most braiding operations (and by extend Cliffords) makes their behavior unintuitive compared to their qubit counterparts, it is worth emphasizing that operations corresponding to Householder reflections with weight 2 act as simple permutation operators. This is further elaborated on in Appendix C.1.
@@ IV Fermionic Stabilizers
A simple example is provided in Appendix C.2, and further such stabilizers can be found in e.g. [9, 15, 12, 10, 11].
@@ IV.1 Symplectic Representation
The reason why a stabilizer can have at most $n$ independent generators becomes apparent when one considers its binary equivalent:
@@
Due to fermion superselection we only care about isotropic subspaces of the even-parity subspace $\mathbb{F}_{2}^{2n,+}$. In this case any maximally isotropic subspace can be written as [^8] Note that the orthogonal matrix $S$ is not at all unique.
@@
where $M_{\text{std}}$ is the canonical maximally isotropic subspace, here provided in terms of a $n\times 2n$ matrix with the basis elements as row vectors:
@@
[…]
A priori, we allow $v$ to have odd parity, as it only affects the leading phases. We will return to this in section IV.2.